Determine the value of k for which the system of linear equations has no solution.
4x - y = -16
16x + ky = 11

Answer by ankor@dixie-net.com(22740)

4x - y = -16
16x + ky = 11
:
Multiply the 1st equation by 4
16x - 4y = -64
16x + ky = 11
It can be seen if k=-4, the system can't have a solution
16x - 4y = -64
16x - 4y = 11